OpenStudy (anonymous):

HELP! A stock solution of 12.5 M HCl is used to make 3.0 L of 1.6 M HCl. Explain (in your own words) in detail how you would make this solution. Show all calculations needed.

OpenStudy (diamondboy):

hi!

OpenStudy (anonymous):

@diamondboy Hi.

OpenStudy (anonymous):

so do you know the dilution formula? \[M _{1}V _{1}=M _{2}V _{2}\]

OpenStudy (anonymous):

on one side you have 12.5M, and the other side is 3L and 1.6M, so solve for the missing volume

OpenStudy (anonymous):

M1V1 = M2V2 M1=12.5 M2 = 1.6 V2 = 3L

OpenStudy (anonymous):

Therefore 12.5 times l = 1.6 times 3

OpenStudy (anonymous):

12.5l = 4.8

OpenStudy (anonymous):

Yup! so now I = 4.8/12.5 = answer and you're done!

OpenStudy (anonymous):

l = 4.8/12.5

OpenStudy (anonymous):

hence ans = 0.384

OpenStudy (anonymous):

looks right